OpenStudy (anonymous):

I \[(x^2)^2\] the same as \[x^4\]

OpenStudy (anonymous):

*Is

OpenStudy (jhannybean):

yes, because (x^2)^2 = x^(2*2)

Parth (parthkohli):

\[(x^2)^2 = x^2 \times x^2 = x\times x \times x \times x = x^4 \]

OpenStudy (jhannybean):

(a^x)^x = a^(x*x)

OpenStudy (jhannybean):

x being any number of course..

OpenStudy (anonymous):

So if i have ((x^4)/-x)^2) is it the same as x^6/-x^2

OpenStudy (jhannybean):

x^4/-x^2 = -x^(4-2)

OpenStudy (unklerhaukus):

\[(x^n)^m=\overbrace{\Big(\overbrace{x\times x\times x\times\dots\times x}^{\text {n times}} \big)\times\Big(\overbrace{x\times x\times x\times\dots\times x}^{\text {n times}} \big)\times\dots\times\Big(\overbrace{x\times x\times x\times\dots\times x}^{\text {n times}} \big)}^{\text{ m times}}\]

OpenStudy (jhannybean):

WHOAH

OpenStudy (anonymous):

But is ((x^4)/-x)^2) = x^6/-x^2 if it dose the over all resolution is then -x^6

OpenStudy (unklerhaukus):

\[\frac{x^n}{x^m}=\frac{\overbrace{x\times x\times x\times\dots\times x}^{\text {n times}}}{\overbrace{x\times x\times x\times\dots\times x}^{\text {m times}}} =\frac{\overbrace{x\times x\times x}^{\text {n-m times}}\times\cancel{\dots\times x}}{{1\times\cancel{x\times x\times x\times\dots\times x}}}={x^{n-m}}{} \]

OpenStudy (anonymous):

Is the simplified version of the -x^6? If so then when -x^6=10^9 dose that = x=-10^3

OpenStudy (unklerhaukus):

i dont understand what you are saying @merh

OpenStudy (anonymous):

Sorry, I am trying to solve this log problem and exponents just have a way with messing with my head so i am trying to clarify things. I started with : \[10^9=((x^2)^2/-x)^2\] and simplified it to: \[10^9=(x^6/-x)^2\] and then \[10^9=-x^6\] and finally \[x=10^3\] I just wanted to make sure that this was the correct answer.

OpenStudy (anonymous):

*\[10^9=(x^4/−x)^2\] the second time not x^6

OpenStudy (unklerhaukus):

ok there is a mistake \[10^9=\left(\frac{x^4}{−x}\right)^2=\big(-x^{4-1}\big)^2=(-x^3)^2=(-x^3)\times(-x^3)=x^6\]

OpenStudy (unklerhaukus):

and for the final step you have to take the sixth root of both sides \[10^9=x^6\\\sqrt[6]{10^9}=\sqrt[6]{x^6}\\10^{9/6}=x\]

OpenStudy (anonymous):

I though that you could subtract exponents for division? Is it not 10^3?

OpenStudy (unklerhaukus):

we are not dividing on this step, we are taking the sixth root

OpenStudy (anonymous):

is the answer 10^3 yes or no?

OpenStudy (unklerhaukus):

nope